# Learnability is a Compact Property

### v1

@@ 3.4 Distribution-family Learning
The analysis of PAC learning with respect to more flexible distribution classes than the realizable and agnostic cases falls largely under the purview of distribution-family learning [11]. Formally, a problem in distribution-family learning of a class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ is defined by a family of distributions $\mathbb{D}$ over $\mathcal{X}$, such that unlabeled datapoints are drawn from a distribution $D\in\mathbb{D}$ and labeled by a hypothesis $h\in\mathcal{H}$ (in the realizable case) or arbitrarily (in the agnostic case).
Notably, distribution-family learning has infamously resisted any characterization of learnability, combinatorial or otherwise, for the 40 years since its inception. In fact, there is some evidence to suggest that no such characterization may exist [31]. Furthermore, it is a setting in which uniform convergence fails to characterize learning, rendering ineffective many of the standard and most celebrated techniques of learning theory.
@@
Definition 3.16 is sufficiently flexible that we may apply it to distribution-family learning with $\mathcal{Z}=\mathcal{X}$ or to PAC learning over arbitrary distribution classes $\mathbb{D}$ with $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$. Though it may appear overly restrictive at first glance, note that well-behavedness is satisfied not only for ordinary PAC learning in the agnostic and realizable cases, but also for learning of partial concept classes in the realizable case [6] and for the EMX learning of [10]. [^2] We remark that the EMX problem is not presented as a supervised learning problem in Ben-David et al. [10], but can easily be seen as a a binary classification problem with $\mathbb{D}$ containing all realizable and discrete distributions. Crucially, well-behavedness permits us to study PAC learning by way of transductive error.
Proposition 3.17 (Informal Proposition B.3).
Let $\mathcal{H}$ be a hypothesis class and $\mathbb{D}$ a well-behaved family of realizable distributions (i.e., $\inf_{\mathcal{H}}L_{D}(h)=0\;\forall D\in\mathbb{D}$). Fix a bounded loss function. Then the transductive and PAC sample complexities of $\mathcal{H}$ differ by at most a logarithmic factor.
It is now immediate from the proof of Theorem 3.6 that distribution-family learning is a setting in which the sample complexity of learning is $\mathcal{H}$ equals the sample complexity of learning its most challenging finite projections. In the following, we let $\mathbb{D}|_{X}\subseteq\mathbb{D}$ denote the distributions of $\mathbb{D}$ which place full measure on $X\subseteq\mathcal{X}$.
@@
Let $\mathcal{X}$ be an arbitrary domain, $\mathcal{Y}$ a proper metric space, and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Let $\mathbb{D}$ be a family of well-behaved, realizable distributions. Then the following are equivalent:
@@
    $\mathcal{H}$ is learnable with respect to $\mathbb{D}$ with transductive sample function $m:(0,1)\to\mathbb{N}$.
@@
    For any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\mathcal{H}^{\prime}$ is learnable with respect to $\mathbb{D}|_{X}$ with transductive sample function $m:(0,1)\to\mathbb{N}$.
Note that Theorem 3.18 holds for the natural definition of transductive learning with respect to $\mathbb{D}$, i.e., in which the adversary must select a sequence of unlabeled datapoints $S\in\mathcal{X}^{n}$ which lie in the support of some $D\in\mathbb{D}$. An immediate consequence of Theorem 3.18 is that the EMX learning of Ben-David et al. [10] is within the reach of compactness. In particular, the learnability of such problems can be detected by examining all their finite projections, provided that no restrictions are placed upon learners. In Ben-David et al. [10], however, learners were required to only emit hypotheses in $\mathcal{H}$. Our work demonstrates that the nature of their undecidability result — in which the learnability of a class $\mathcal{H}$ is determined entirely by its cardinality, despite all of its projections being easily learned — could otherwise not appear in supervised learning with metric losses.
Acknowledgements
Julian Asilis was supported by the USC Viterbi School of Engineering Graduate School Fellowship and by the Simons Foundation. Siddartha Devic was supported by the Department of Defense through the National Defense Science & Engineering Graduate (NDSEG) Fellowship Program. Shaddin Dughmi was supported by NSF Grant CCF-2009060. Vatsal Sharan was supported by NSF CAREER Award CCF-2239265 and an Amazon Research Award. Shang-Hua Teng was supported in part by NSF Grant CCF-2308744 and the Simons Investigator Award from the Simons Foundation. We thank Cameron Freer and Nate Ackerman for help in designing Figure 1.
@@ Appendix A Proper metric spaces
Section 2.3 restricts focus to the case of supervised learning with respect to proper loss functions. Let us expand briefly upon this condition, and present an equivalent definition.
@@
Recall now that $\mathbb{R}^{n}$ endowed with the usual Euclidean norm is a proper metric space, owing to the Heine-Borel theorem. Invoking the equivalence of all norms on $\mathbb{R}^{n}$, it follows that $\mathbb{R}^{n}$ endowed with any norm enjoys the structure of a proper metric space.
@@
    $\mathbb{R}^{n}$, with any norm.
@@
    Any closed subset of $\mathbb{R}^{n}$.
Regarding necessary conditions for properness, note that all proper metric spaces are complete. Thus subsets of $\mathbb{R}^{n}$ which are not closed will not be proper, e.g., $\mathbb{Q}\subseteq\mathbb{R}^{1}$. See, e.g., [41] for additional discussion and properties of proper metric spaces, which are sometimes referred to as Heine-Borel metric spaces.
@@ Appendix B Omitted proofs
Consider now the poset $\mathcal{P}$ whose elements are finitely satisfiable partial assignments of $(\epsilon+\delta)$-apportionments to $R$. The partial ordering is such that $\phi_{1}\leq\phi_{2}$ if $\phi_{2}$ agrees with all assignments of apportionments made by $\phi_{1}$ and perhaps assigns additional apportionments. It is straightforward to see that chains in $\mathcal{P}$ have upper bounds: let $\mathcal{C}\subseteq\mathcal{P}$ be a chain and define $\phi_{\mathcal{C}}$ to be the “union” of assignments in $\mathcal{C}$, i.e., $\phi_{\mathcal{C}}$ leaves $r\in\mathbb{R}$ unassigned if all $\phi\in\mathcal{C}$ leave $r$ unassigned, otherwise it assigns $r$ to the unique apportionment used by the assignments in $\mathcal{C}$.
@@
Upon assigning each such variable $\ell_{i}$ to an element of $\mathcal{Y}$ (semantically, a completion of its “?” entry corresponding to a query at test time) the function $r$ outputs the value
@@
First note that $G$ is finitely $R$-matchable if and only if P. Hall’s condition holds for all finite subsets $R^{\prime}$ of $R$. That is, if and only if $|N(R^{\prime})|\geq|R^{\prime}|$, where $N(R^{\prime})$ denotes the set of neighbors $R^{\prime}$ has in $L$, by [23]. We refer to a finite subset $R^{\prime}$ of $R$ as a blocking set if $|N(R^{\prime})|=|R^{\prime}|$. We say a node $\ell\in L$ is contained in a blocking set if there exists a blocking set $R^{\prime}$ such that $\ell\in N(R^{\prime})$.
@@
B.4 Proof of Proposition 3.17
Proposition B.3.
Let $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ be a hypothesis class and $\mathbb{D}$ a well-behaved family of distributions which are realizable (i.e., $\inf_{\mathcal{H}}L_{D}(h)=0\;\forall D\in\mathbb{D}$). Fix a loss function taking values in $[0,1]$. Then the following inequality holds for all $\epsilon\in(0,1)$ and an absolute constant $c>0$:
$m_{\operatorname{Trans},\mathcal{H}}(c\cdot\epsilon)\leq m_{\operatorname{PAC},\mathcal{H}}(\epsilon,\epsilon)\leq m_{\operatorname{Trans},\mathcal{H}}(\epsilon/c)\cdot c\cdot\log(1/\epsilon).$
Proof sketch.
The proof is nearly identical to that of [7, Proposition 3.6]. In particular, let $m_{\operatorname{Exp},\mathcal{H}}$ denote the sample complexity of learning $\mathcal{H}$ in the expected regime (i.e., $m_{\operatorname{Exp},\mathcal{H}}(\epsilon)$ equals the number of datapoints needed to incur expected error at most $\epsilon$). Then for an absolute constant $c$, we have
$m_{\operatorname{Exp},\mathcal{H}}(\epsilon)\leq m_{\operatorname{PAC},\mathcal{H}}(\epsilon/2,\epsilon/2)\leq m_{\operatorname{Exp},\mathcal{H}}(\epsilon/c)\cdot c\cdot\log(1/\epsilon).$
The first inequality follows immediately from the fact that the loss function is bounded above by 1. The second inequality follows from a repetition argument, i.e., a learner attaining expected error $\leq\epsilon$ on samples of size $k$ can be boosted to attain expected error $\leq\frac{2}{\epsilon}$ with probability $\geq 1-\epsilon$ by using an additional factor of $O(\log(1/\epsilon))$ many samples.
The first inequality follows inequality from a standard leave-one-out argument of [29]. The second inequality follows the fact that for any $S=((x_{1},y_{1}),\ldots,(x_{n},y_{n}))$, there exists an $m_{n}\in\mathbb{N}$ such that $m_{n}$ many independent draws from the uniform distribution over the entries of $S$ has probability at least $\frac{1}{\epsilon}$ of containing exactly $n-1$ elements of $S$, as detailed in [7, Lemma A.1]. Crucially, the uniform distribution over $S$ lies in $\mathbb{D}$ owing to well-behavedness of $\mathbb{D}$. The claim follows from both of the established chains of inequalities. ∎

### v2

@@ 3.5 Distribution-family Learning
The analysis of PAC learning with respect to more flexible distribution classes than the realizable and agnostic cases falls largely under the purview of distribution-family learning (Benedek and Itai, 1991). Formally, a problem in distribution-family learning of a class $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ is defined by a family of distributions $\mathbb{D}$ over $\mathcal{X}$, such that unlabeled datapoints are drawn from a distribution $D\in\mathbb{D}$ and labeled by a hypothesis $h\in\mathcal{H}$ (in the realizable case) or arbitrarily (in the agnostic case).
Notably, distribution-family learning has infamously resisted any characterization of learnability, combinatorial or otherwise, for the 40 years since its inception. In fact, there is some evidence to suggest that no such characterization may exist (Lechner and Ben-David, 2023). Furthermore, it is a setting in which uniform convergence fails to characterize learning, rendering ineffective many of the standard and most celebrated techniques of learning theory.
@@
Definition 3.21 is sufficiently flexible that we may apply it to distribution-family learning with $\mathcal{Z}=\mathcal{X}$ or to PAC learning over arbitrary distribution classes $\mathbb{D}$ with $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$. Though it may appear overly restrictive at first glance, note that well-behavedness is satisfied not only for ordinary PAC learning in the agnostic and realizable cases, but also for learning of partial concept classes in the realizable case (Alon et al., 2022) and for the EMX learning of Ben-David et al. (2019). In particular, though EMX learning is not presented as a supervised learning problem in Ben-David et al. (2019), it can be seen as a binary classification problem over a domain $\mathcal{X}$ for which $\mathcal{H}$ consists of those functions outputting finitely many 1’s and $\mathbb{D}$ contains all realizable, discrete distributions placing all $\mathcal{Y}$-mass on the label 1.
Crucially, well-behavedness permits us to study PAC learning by way of transductive error.
Proposition 3.22.
Let $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ be a hypothesis class and $\mathbb{D}$ a well-behaved family of distributions which are realizable (i.e., $\inf_{\mathcal{H}}L_{D}(h)=0\;\forall D\in\mathbb{D}$). Fix a loss function taking values in $[0,1]$. Then the following inequality holds for all $\epsilon,\delta\in(0,1)$ and the constant $e\approx 2.718$:
$m_{\operatorname{Trans},\mathcal{H}}(e\cdot(\epsilon+\delta))\leq m_{\operatorname{PAC},\mathcal{H}}(\epsilon,\delta)\leq O\big{(}m_{\operatorname{Trans},\mathcal{H}}(\epsilon/2)\cdot\log(1/\delta)\big{)}.$
Proof.
The argument follows identically as in Lemma 3.19. Crucially, $m_{\operatorname{Trans},\mathcal{H}}(\epsilon)\leq m_{\operatorname{Exp},\mathcal{H}}(\epsilon/e)$ due to the fact that a learner attaining low expected error must perform well on the uniform distribution over any particular learning instance $S=(x_{i},y_{i})_{i\in[n]}$ — owing to well-behavedness of $\mathbb{D}$ — and can thus be used to create a good transductive learner, as described in Lemma 3.19. ∎
Note that Proposition 3.22 holds for the natural definition of transductive learning with respect to $\mathbb{D}$, i.e., in which the adversary must select a sequence of unlabeled datapoints $S\in\mathcal{X}^{n}$ which lie in the support of some $D\in\mathbb{D}$. It is now immediate from the proof of Theorem 3.6 that distribution-family learning is a setting in which the transductive sample complexity of learning is $\mathcal{H}$ equals the sample complexity of learning its most challenging finite projections. In the following, we let $\mathbb{D}|_{X}\subseteq\mathbb{D}$ denote the distributions of $\mathbb{D}$ which place full measure on $X\subseteq\mathcal{X}$.
@@
Let $\mathcal{X}$ be an arbitrary domain, $\mathcal{Y}$ a proper metric space, and $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$ a hypothesis class. Let $\mathbb{D}$ be a family of well-behaved, realizable distributions. Then the following are equivalent for any $m:\mathbb{R}_{>0}\to\mathbb{N}$:
@@
    $\mathcal{H}$ is learnable with respect to $\mathbb{D}$ with transductive sample function $m$.
@@
    For any finite $X\subseteq\mathcal{X}$ and finite $\mathcal{H}^{\prime}\subseteq\mathcal{H}|_{X}$, $\mathcal{H}^{\prime}$ is learnable with respect to $\mathbb{D}|_{X}$ with transductive sample function $m$.
As in Section 3.4, Proposition 3.22 applied to Theorem 3.23 immediately yields an almost-exact form of compactness for distribution-family PAC learning with realizable and well-behaved distribution classes. This demonstrates that the learnability of even problems as exotic as EMX learning can be detected by examining all their finite projections, provided that no restrictions are placed upon learners. In Ben-David et al. (2019), however, learners were required to only emit hypotheses in $\mathcal{H}$. Our work demonstrates that the nature of their undecidability result — in which the learnability of a class $\mathcal{H}$ is determined entirely by its cardinality, despite all its projections being easily learned — could otherwise not appear in supervised learning with metric losses.
4 Conclusion
In this work, we studied the conditions under which the sample complexity of learning a class $\mathcal{H}$ can be detected by examining its finite projections. Notably, we established exact compactness results for transductive learning with a broad class of proper or continuous loss functions, across both realizable and agnostic learning. Using bounds relating the transductive and PAC models, we were able to transfer many of our results (in an approximate form) to realizable PAC learning. We leave as an open problem whether compactness of agnostic transductive sample complexities can fail by more than a factor of 2 for arbitrary (improper) metric losses. Additional future work includes better understanding the relationship between the transductive and PAC models in the agnostic case, and examining compactness for loss functions which do not satisfy any of our properness, metric, or continuity conditions (though they may be of somewhat limited interest in learning theory). It would also be of interest to study the compactness of error rates in settings other than supervised learning, such as online or unsupervised learning.
Acknowledgments and Disclosure of Funding
Julian Asilis was supported by the Simons Foundation and by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-1842487. Siddartha Devic was supported by the Department of Defense through the National Defense Science & Engineering Graduate (NDSEG) Fellowship Program. Shaddin Dughmi was supported by NSF Grant CCF-2009060. Vatsal Sharan was supported by NSF CAREER Award CCF-2239265 and an Amazon Research Award. Shang-Hua Teng was supported in part by NSF Grant CCF-2308744 and the Simons Investigator Award from the Simons Foundation. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of any of the sponsors such as the NSF.
@@ Appendix A Proper metric spaces
Several of our results concern proper metric spaces. Let us expand briefly upon this condition, and present an equivalent definition.
@@
Recall now that $\mathbb{R}^{d}$ endowed with the usual Euclidean norm is a proper metric space, owing to the Heine-Borel theorem. Invoking the equivalence of all norms on $\mathbb{R}^{d}$, it follows that $\mathbb{R}^{d}$ endowed with any norm enjoys the structure of a proper metric space.
@@
    $\mathbb{R}^{d}$, with any norm.
@@
    Any closed subset of $\mathbb{R}^{d}$, with any norm.
Regarding necessary conditions for properness, note that all proper metric spaces are complete. Thus subsets of $\mathbb{R}^{d}$ which are not closed will not be proper, e.g., $\mathbb{Q}\subseteq\mathbb{R}^{1}$. See, e.g., Williamson and Janos [1987] for additional discussion and properties of proper metric spaces, which are sometimes referred to as Heine-Borel metric spaces.
@@ Appendix B Omitted proofs
Consider now the poset $\mathcal{P}$ whose elements are finitely satisfiable partial assignments of $(\epsilon+\delta)$-apportionments to $R$. The partial ordering is such that $\phi_{1}\leq\phi_{2}$ if $\phi_{2}$ agrees with all assignments of apportionments made by $\phi_{1}$ and perhaps assigns additional apportionments. It is straightforward to see that chains in $\mathcal{P}$ have upper bounds: let $\mathcal{C}\subseteq\mathcal{P}$ be a chain and define $\phi_{\mathcal{C}}$ to be the “union" of assignments in $\mathcal{C}$, i.e., $\phi_{\mathcal{C}}$ leaves $r\in\mathbb{R}$ unassigned if all $\phi\in\mathcal{C}$ leave $r$ unassigned, otherwise it assigns $r$ to the unique apportionment used by the assignments in $\mathcal{C}$.
@@
Upon assigning each such variable $\ell_{i}$ to an element of $\mathcal{Y}$ (semantically, a completion of its “?" entry corresponding to a query at test time) the function $r$ outputs the value
@@
First note that $G$ is finitely $R$-matchable if and only if P. Hall’s condition holds for all finite subsets $R^{\prime}$ of $R$. That is, if and only if $|N(R^{\prime})|\geq|R^{\prime}|$, where $N(R^{\prime})$ denotes the set of neighbors $R^{\prime}$ has in $L$, by Hall [1935]. We refer to a finite subset $R^{\prime}$ of $R$ as a blocking set if $|N(R^{\prime})|=|R^{\prime}|$. We say a node $\ell\in L$ is contained in a blocking set if there exists a blocking set $R^{\prime}$ such that $\ell\in N(R^{\prime})$.